# Demystifying the Compression of Mixture-of-Experts Through a Unified Framework

### v2

@@ 5.1 Expert Slimming
Table 2: Performance of Pruning on MoE. We consider two mainstream pruning methods (i.e., Wanda [4] and SparseGPT [45]) under $50\%$ unstructured sparsity and $2$:$4$ semi-structured sparsity.
@@
Pruning Maintains Performance but Faces Deployment Limitations.
In Table 2, we evaluate representative pruning algorithms (i.e., Wanda [4], SparseGPT [45]) on Mixtral-8$\times$7B and DeepSeek-MoE-16B. Since DeepSeek-MoE-16B utilizes both shared experts and normal experts, we conduct an ablation study on whether to prune shared experts, as discussed in Appendix D. We find that unstructured pruning preserves more than $95\%$ of performance. However, it is not compatible with existing hardware. Conversely, the hardware-friendly semi-structured pruning (i.e., $4$:$8$ and $2$:$4$ patterns) undergoes a significant performance drop. Nevertheless, according to Lu et al. [11], semi-structured sparsity is ineffective in speeding up MoE models.
Table 3: Performance of Quantization on MoE. We utilize GPTQ [5] and AWQ [42] as the quantization methods for 4-bit compression.
@@
Quantization Demonstrates Promising Performance and Efficiency.
In Table 3, we evaluate the impact of 4-bit quantization on MoE. Quantization offers two major benefits: it maintains the comparable performance of the original models and significantly reduces memory costs. Specifically, the quantized models achieve over $98\%$ of the original performance while using less than $30\%$ of the memory. Moreover, when quantized with AWQ [42], Mixtral-8$\times$7B and DeepSeek-MoE-16B achieve impressive speedups of ×$\times 5.08$ and ×$\times 3.16$, respectively. This demonstrates that 4-bit quantization is an effective technique for deploying MoE models in resource-constrained environments.
5.2 Expert Trimming
Figure 3: Evaluation of Expert Drop. We consider two strategies: layer-wise (dotted lines) and global (solid lines).
Expert Drop Degrades Performance and Offers Limited Efficient Gains.
While experts are specific structures in MoE, not all experts hold equal significance. Figure 9 visualizes the distribution of expert-wise importance scores, highlighting this variability. To systematically drop experts at varying proportions, we conduct experiments using both layer-wise and global dropping approaches (see Appendix A.3). Given the importance of shared experts in DeepSeek-MoE-16B, we only dropped normal experts. Under both settings, Expert Drop causes consistent performance degradation. For example, dropping $25\%$ of experts in Mixtral-8$\times$7B results in a $23\%$ performance drop on the MMLU task.
The efficiency improvement from Expert Drop is also marginal. For instance, dropping $12.5\%$ of experts results in less than a $1\%$ speedup, despite significant performance losses. More results are available in Appendix E.
Layer Drop Preserves the Performance Efficiently.
To verify the feasibility of Layer Drop, we visualize feature similarity across different modules in Figure 4. This visualization shows a high level of similarity between the inputs of the MoE layer normalization module (Norm) and the outputs of the MoE layer. In contrast, the low similarity between the inputs and outputs of MoE layers indicates the infeasibility of removing only the MoE layers. Results from Figure 5 show that Layer Drop preserves performance within a wide range of compression ratio, e.g. $1\%$ performance drop on MMLU when dropping 8 layers for Mixtral-8$\times$7B, revealing significant redundancy in MoE layers.
Block Drop Further Promotes Efficiency.
While Layer Drop maintains the performance of the original models, it still retains the computation-costly attention layers. To address this, we extend the approach by dropping larger structures, specifically blocks that include both FFN and attention layers. We first visualized the similarity for different blocks in Figure 8, where both Mixtral-8$\times$7B and DeepSeek-MoE-16B demonstrate high similarity between specific blocks. Based on this, we conduct the empirical study by varying the number of dropped blocks. Surprisingly, as shown in Figure 6, even after removing 4 blocks, the models maintained over $90\%$ of the original performance. Since Block Drop removes computationally expensive attention layers, it outperforms Layer Drop by a large margin in terms of both memory and inference cost, as illustrated in Figure 8. By focusing on higher-level structures, Block Drop and Layer Drop provide significant efficiency improvements while maintaining acceptable performance levels.
5.3 Integration of Expert Slimming and Expert Trimming
We further investigate the integration of Expert Slimming and Expert Trimming techniques. Given the superior average performance and practical efficiency of quantization, we use it for Expert Slimming. For Expert Trimming, we include all three methods to offer a comprehensive comparison.
Quantization Preserves the Performance of Expert Trimming. As shown in Table 4, the integration of Expert Slimming and Expert Trimming significantly enhances overall efficiency. Quantization can be seamlessly combined with three different levels of dropping, achieving comparable performance. For instance, after quantization, the average performance of Layer Drop and Block Drop is nearly the same, maintaining more than $90\%$ of the performance of the original models.
The Integration Significantly Promotes Efficiency. For efficiency considerations, we measure speedup, FLOPs, and memory usage. Since the original Mixtral-8$\times$7B cannot be deployed on a single Nvidia A100 GPU, we utilize two GPUs for all models to ensure a fair comparison. In Table 4, quantization promotes efficiency by a large margin. Different Expert Trimming strategies showcase different advantages. Specifically, Expert Drop contributes to reducing memory usage but its speedup is marginal. Layer Drop and Block Drop excel in speedup as illustrated in Figure 8, with Block Drop demonstrating both higher performance and greater speedup. Considering all settings, the combination of Block Drop and quantization offers the best efficiency with comparable performance: a 6.05×6.05$\times$6.05 × speedup with only $20.0$GB memory usage, while maintaining over $92\%$ of the performance on Mixtral-8$\times$7B, making it available to be deployed on a NVIDIA RTX 3090 GPU.
@@ 5.3 Integration of Expert Slimming and Expert Trimming
+ E2/8 1.02×1.02$\times$1.02 × 54.4T 66.7GB 53.2 77.7 80.5 52.2 46.2 81.7 55.6 76.8 65.5
@@
[…]
DeepSeek-MoE-16B
Baseline 11.7T 30.8GB 48.1 72.4 77.3 37.9 44.0 80.4 63.9 70.3 61.8
[…]
+ E16/64 1.03×1.03$\times$1.03 × 11.7T 23.9GB 45.0 67.1 75.6 31.8 42.2 80.2 59.9 70.0 59.0
[…]
6 Discussion
With the advent of open-source LLMs incorporated with MoE (e.g., Mixtral [7]), the machine learning community has rallied around the philosophy of utilizing MoE to scale large language models and making MoE models more accessible to real-life applications. Our research delves into making MoE models more efficient and accessible, bringing forth significant discoveries and future pathways:
Expert Slimming Exhibits Superior Compression Performance. Anthough MoE scales models with multiple experts, it causes excessive memory usage and slows down the inference. Expert Slimming mitigates this by directly compressing individual experts, significantly promoting inference speed and memory.
Expert Trimming Further Creates Efficient Architectures. Scaling with MoE architectures also introduces redundant structures that showcase highly similar hidden features. Expert Trimming removes unimportant modules and creates more efficient architectures.
Model/Task-Specific Compression. Mixtral and DeepSeek-MoE exhibit different properties of redundancy: compared to Mixtral, DeepSeek-MoE has higher expert-redundancy (more experts with low importance scores in Figure 9) and lower hidden feature similarities at both layer and block levels, which is shown in Figure 4 and 8. On the other hand, MoE models can be compressed by varying ratios depending on the task. Establishing an under-explored correlation between the specific task and the corresponding optimal compression ratio.
Post-Compression Training. In this work, we compress MoE models in a training-free manner. However, our framework facilitates the acquisition of compressed, efficient MoE architectures that can be continuously trained to enhance performance [4, 59, 21, 30]. This approach enables MoE architectures to be improved in both efficiency and performance.
7 Conclusion
In this paper, we introduced a unified framework of MoE compression, facilitating to systematically understand the efficiency issue of MoE and identify the new design space to further improve the performance. Based on this framework, we propose a comprehensive recipe that integrates Expert Slimming and Expert Trimming to further enhance efficiency. Our proposed methods and insights not only address current challenges but also set the stage for future advancements in the field of MoE.
Appendix A Implementation Details
A.1 Models and Datasets
Models.
For our experiments, we employed Mixtral-8$\times$7B [7] and DeepSeek-MoE-16B [8]. Mixtral-8$\times$7B utilizes 8 experts for MoE layers and activates the top two for each input token. In contrast, DeepSeek-MoE-16B employs an MLP in the first block and utilizes two shared experts with additional 64 experts within MoE layers in other blocks.
Datasets.
For compression experiments, we used the C4 dataset [60], with 128 samples and an input sequence length of 2,048, following the setup in [4, 11, 42, 5]. To evaluate model performance, we report normalized zero-shot accuracy on the LM-harness benchmark, which includes multiple tasks: ARC-C [61], BoolQ [62], HellaSwag [63], MMLU [64], OBQA [65], PIQA [66], RTE [67], and WinoGrande [68]. The evaluation code is based on EleutherAI LM Harness [69].
A.2 Implementation Details of Expert Slimming
Both Expert Slimming methods (i.e., pruning and quantization) require calibration data to estimate input statistics. To control this variable, we use $128$ samples from the C4 dataset [60] as the calibration dataset for pruning. For quantization, we follow the default settings of GPTQ  [^1] https://github.com/AutoGPTQ/AutoGPTQ and AWQ  [^2] https://github.com/casper-hansen/AutoAWQ, using $128$ random samples from Alpaca [70] and Pile [71], respectively. We use the default group size $128$ for Mixtral-8$\times$7B and $64$ for DeepSeek-MoE-16B.
A.3 Implementation Details of Expert Drop
The Expert Drop compresses MoE by preserving only important experts $\{\boldsymbol{E}_{i}\}_{i\in\mathcal{T}^{\prime}}$ while removing others, where $\mathcal{T}^{\prime}$ is determined by the importance scores $\{\boldsymbol{S}(\boldsymbol{E}_{i})\}_{i\in\mathcal{T}}$. Following Muzio et al. [21], we measure the importance scores through the averaged routing scores of a batched data $\mathcal{X}$, i.e., $\{\boldsymbol{S}(\boldsymbol{E}_{i})\}=\frac{1}{|\mathcal{X}|}\sum_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{G}_{i}(\boldsymbol{x})$, and consider two dropping strategies for Expert Drop: layer-wise dropping and global dropping.
$\mathcal{T}^{\prime(l)}=\{\boldsymbol{E}_{t}^{(l)}\},\quad\text{where}\quad\boldsymbol{S}(\boldsymbol{E}_{t}^{(l)})\in\mathrm{TopK}(\{\boldsymbol{S}(\boldsymbol{E}_{i}^{(l)})\}_{i=1}^{n},n^{\prime}).$ (10)
$\mathcal{T}^{\prime(l)}=\{\boldsymbol{E}_{t}^{(l)}\},\quad\text{where}\quad\boldsymbol{S}(\boldsymbol{E}_{t}^{(l)})\in\mathrm{TopK}\Big{(}\bigcup_{j=1}^{m}\{\boldsymbol{S}(\boldsymbol{E}_{i}^{(j)})\}_{i=1}^{n},n^{\prime}L\Big{)}.$ (11)
Appendix B Analysis on the Dropping Patterns
In this section, we make further analysis on the dropping patterns of Expert Trimming methods.
B.1 Expert Drop
Figure 9: Distribution of Normalized Importance Scores $\boldsymbol{S}$ for Expert Drop. We highlight the density of scores under different drop ratios with different colors.
Score Distribution Directs Expert Drop.
The distribution of importance scores is informative to determine the proportion of dropped experts. In Figure 9, we visualize the score distribution of Expert Drop for Mixtral-8$\times$7B and DeepSeek-MoE-16B, respectively. DeepSeek-MoE-16B, which allocates more experts, shows a left-skewed distribution where more experts have low importance scores. In contrast, Mixtral-8$\times$7B demonstrates a right-skewed distribution, with only a few experts being deemed unimportant. This significantly different distribution results in different resistance capability against Expert Drop, where DeepSeek-MoE-16B can drop much more experts than Mixtral-8$\times$7B while maintaining competitive performance, as demonstrated in Table 4 and Figure 13.
Figure 10: Distribution of Dropped Experts for Expert Drop. We visualize of the dropped experts under different drop ratios, where the dropped experts are colored from yellow to blue as the drop ratio increases.
Global Expert Drop Removes Experts Fine-Grainedly.
B.2 Layer Drop & Block Drop
(a) Layer Drop
(b) Block Drop
Figure 11 visualizes the remaining and dropped layers/blocks as the number of dropped modules increases. Both MoE architectures exhibit similar patterns in Layer Drop and Block Drop: initially, both models tend to drop the deeper layers, followed by the shallower ones. These findings are consistent with Xu et al. [72], which suggests that deeper layers tend to be more redundant.
Appendix C Ablation Study on the Data Choices
The routing scores in MoE models exhibit a significant correlation with the input data, as evidenced by [30, 8, 11]. This relationship underscores the importance of data selection in the application of Expert Drop, which assesses the importance of experts by examining their routing scores. Similarly, the same question holds for Layer Drop and Block Drop: Does data selection affect the feature similarity across different layers? To address this, we performed ablation studies on Mixtral-8$\times$7B, scrutinizing both the number of samples and the dataset types used for feature extraction, to determine how data choice impacts decisions to drop layers or blocks. Results are shown in Figure 12.
(a) Similarities with different number of samples.
(b) Similarities with different datasets.
Figure 12: Influence of Data Choices on Feature Similarity. We measure the similarity among layers and blocks on Mixtral-8$\times$7B. (a) The normalized similarity calculated using different number of samples from C4 [60]. (b) The normalized similarity calculated using $1,024$ samples from different datasets, i.e., C4, Lima [73] and MetaMathQA [74].
Layer Drop and Block Drop are Data Robust.
In Figure 12(a), we note that feature similarity remains relatively stable across different layers as the sample size increases, indicating that Layer Drop and Block Drop maintain consistency regardless of sample quantity. This confirms that using $128$ samples suffices for computing similarity, which is adopted for all our experiments. Similarly, Figure 12(b) shows that varying the datasets, from pretraining with C4 to instruction tuning with Lima and MetaMathQA, does not significantly alter feature similarity. This demonstrates the resilience of Layer Drop and Block Drop to variations in data distribution.
Appendix D Ablation Study on Shared Experts in DeepSeek-MoE-16B
@@
While most MoE models follow Equation 2 to implement the experts, models like DeepSeek-MoE-16B adopt a residual [33] form of experts, which brings a special scenario to discuss. In the residual MoE, an extra set of $m$ shared experts $\left\{\bar{\boldsymbol{E}}_{1},\bar{\boldsymbol{E}}_{2},\dots,\bar{\boldsymbol{E}}_{m}\right\}$ are always selected by the router $\boldsymbol{G}$ and activated for all inputs. Given an input $\boldsymbol{x}$, the output can be represented as a degenerated form of Equation 2, where the scores of shared experts are fixed to $1$:
$\boldsymbol{y}=\sum\nolimits_{i\in\mathcal{K}}{\boldsymbol{G}(\boldsymbol{x})_{i}\cdot\boldsymbol{E}_{i}(\boldsymbol{x})}+\sum\nolimits_{j=1}^{m}{\bar{\boldsymbol{E}}_{j}(\boldsymbol{x})}.$ (12)
@@
This special form of expert routing may bring a difference in the redundancy distribution of MoE. Here we discuss the influence of shared experts through pruning and present the results in Table 5. We find that pruning without the shared experts will boost the performance at a considerable scale, i.e., $+3.6\%$ and $+1.5\%$ of the averaged accuracy for unstructured pruning with Wanda and SparseGPT, respectively. This finding reveals a different pattern of the inner redundancy in that the shared experts are less compressible compared to the others in residual MoE models, which may inform future work.
@@
Table 5: Ablation Study of Pruning Shared Experts on DeepSeek-MoE-16B. We consider two scenarios, i.e., pruning both shared experts and normal experts (“w/Pruning Shared Experts”) and pruning normal experts only (“w/o Pruning Shared Experts”). We use two mainstream pruning methods (i.e., Wanda [4] and SparseGPT [45]) under both unstructured sparsity ($50\%$) and semi-structured sparsity ($2$:$4$).
@@
Appendix E Full Experimental Results

### v3

@@ Appendix B Expert Slimming
Table 5: Performance of Pruning on MoE. We consider two mainstream pruning methods (i.e., Wanda (Sun et al., 2023) and SparseGPT (Frantar & Alistarh, 2023)) under $50\%$ unstructured sparsity and $2$:$4$ semi-structured sparsity.
@@
Quantization: Better Performance and Greater Efficiency
In Table 6, we evaluate the impact of 4-bit quantization on MoE. The quantization offers two major benefits: it maintains the comparable performance of the original models and significantly reduces memory costs. Specifically, the quantized models achieve over $98\%$ of the original performance while using less than $30\%$ of the memory. When quantized with AWQ (Lin et al., 2024), Mixtral-8$\times$7B and DeepSeek-MoE-16B achieve impressive speedups of ×$\times 5.08$ and ×$\times 3.16$, respectively. This demonstrates that 4-bit quantization is effective for deploying MoE models in resource-constrained environments.
Table 6: Performance of Quantization on MoE. We utilize GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2024) as the quantization methods for 4-bit compression.
@@
Appendix C Analysis on the Dropping Patterns of Expert Drop
Figure 11: Distribution of Normalized Importance Scores $\bm{S}$ for Expert Drop. We highlight the density of scores under different drop ratios with different colors.
Score Distribution Directs Expert Drop
The distribution of importance scores is informative to determine the proportion of dropped experts. In Figure 11, we visualize the score distribution of Expert Drop for Mixtral-8$\times$7B and DeepSeek-MoE-16B. DeepSeek-MoE-16B, which allocates more experts, shows a left-skewed distribution where most experts have low scores. In contrast, Mixtral-8$\times$7B demonstrates a right-skewed distribution, with only a few experts being deemed unimportant. This distribution difference results in different resistance capabilities against Expert Drop, where DeepSeek-MoE-16B can drop more experts than Mixtral-8$\times$7B while maintaining competitive performance, as demonstrated in Table 3 and Figure 13.
Figure 12: Distribution of Dropped Experts for Expert Drop. We visualize the dropped experts under different drop ratios, where the dropped experts are colored from yellow to blue as the drop ratio increases.
Global Expert Drop Removes Experts Fine-Grainedly
Appendix D Ablation Study on Compression Orders
In Section 7, we discussed the combination of Expert Trimming and Expert Slimming. Here we ablate on the orders of compression when combining these two techniques. Results in Table 7 show that the order of Expert Trimming and Expert Slimming doesn’t have a significant influence on the performance, where applying Expert Slimming then Expert Trimming (“S+T”) performs slightly better for Mixtral-8$\times$7B (e.g. +0.5, +0.4 and +0.1 for Expert Drop, Layer Drop and Block Drop, respectively). To this end, we choose “S+T” as the final implementation in our experiments.
Table 7: Ablation results on different orders of Expert Slimming and Expert Trimming. “S+T” denotes first applying Expert Slimming then Expert Trimming, and “T+S” denotes the reversed order.
Mixtral-8×7B
Method ARC-C BoolQ HellaSwag MMLU OBQA PIQA RTE WinoGrande Avg.
Baseline 59.4 84.2 84.0 67.9 46.8 83.8 70.4 75.6 71.5
+ E2/8, AWQ (S+T) 50.7 79.1 78.9 52.4 44.2 81.2 55.6 75.9 64.8
+ E2/8, AWQ (T+S) 50.8 79.9 78.7 49.2 44.4 80.9 55.2 75.4 64.3
@@ Appendix D Ablation Study on Compression Orders
+ B5/32, AWQ (T+S) 50.3 84.7 77.4 65.8 42.0 78.8 70.4 74.0 67.9
DeepSeek-MoE-16B
Method ARC-C BoolQ HellaSwag MMLU OBQA PIQA RTE WinoGrande Avg.
Baseline 48.1 72.4 77.3 37.9 44.0 80.4 63.9 70.3 61.8
+ E16/64, AWQ (S+T) 44.0 66.0 74.5 27.9 42.6 78.5 56.3 67.3 57.1
+ E16/64, AWQ (T+S) 44.7 64.1 74.0 29.0 42.6 79.9 54.2 68.4 57.1
@@
+ B4/28, AWQ (T+S) 41.6 69.4 69.1 35.8 38.6 76.2 50.9 67.0 56.1
Appendix E Ablation Study on Shared Experts in DeepSeek-MoE-16B
@@
Although most MoE models follow Equation 2 to implement the experts, models like DeepSeek-MoE-16B adopt a residual (Rajbhandari et al., 2022) form of experts, which brings a special scenario to discuss. In the residual MoE, an additional set of $m$ shared experts $\left\{\bar{\bm{E}}_{1},\bar{\bm{E}}_{2},\dots,\bar{\bm{E}}_{m}\right\}$ is always selected by the router $\bm{G}$ and activated for all inputs. Given an input $\bm{x}$, the output can be represented as a degenerated form of Equation 2, where the scores of shared experts are fixed to $1$:
$\bm{y}=\sum\nolimits_{i\in\mathcal{K}}{\bm{G}(\bm{x})_{i}\cdot\bm{E}_{i}(\bm{x})}+\sum\nolimits_{j=1}^{m}{\bar{\bm{E}}_{j}(\bm{x})}.$ (12)
@@
This special form of expert routing may bring a difference in the redundancy distribution of MoE. Here we discuss the influence of shared experts through pruning and present the results in Table 8. We find that pruning without the shared experts will boost the performance at a considerable scale, i.e., $+3.6\%$ and $+1.5\%$ of the averaged accuracy for unstructured pruning with Wanda and SparseGPT, respectively. This finding reveals a different pattern of inner redundancy in which the shared experts are less compressible compared to the others in residual MoE models, which may inform future work.
@@
Table 8: Ablation Study of Pruning Shared Experts on DeepSeek-MoE-16B. We consider two scenarios, i.e., pruning both shared experts and normal experts (“w/Pruning Shared Experts”) and pruning normal experts only (“w/o Pruning Shared Experts”). We use two mainstream pruning methods (i.e., Wanda (Sun et al., 2023) and SparseGPT (Frantar & Alistarh, 2023)) under both unstructured sparsity ($50\%$) and semi-structured sparsity ($2$:$4$).
@@
Appendix F Full Experimental Results